OpenStudy (anonymous):

integrate 9dx/1+9x^2

zepdrix (zepdrix):

\[\Large \int\limits\limits\frac{9dx}{1+9x^2}\qquad=\qquad\int\limits\limits\frac{9dx}{1+(3x)^2}\] We want to make the substitution, \(\large 3x=\tan\theta\) Have you done much work with Trig subs yet?

OpenStudy (anonymous):

Yes

zepdrix (zepdrix):

\[\Large dx=\frac{1}{3}\sec^2\theta d \theta\]Applying the substitution gives us,\[\Large \int\limits\frac{9\left(\frac{1}{3}\sec^2\theta \;d \theta\right)}{1+\tan^2\theta}\]Understand how to simplify it down from there? :o

OpenStudy (anonymous):

Do you use arctan?

zepdrix (zepdrix):

After we integrate, yes arctan will be involved near the very end :)

zepdrix (zepdrix):

But for now, it's just some simplification and then cancelling stuff out before we do an easy integration step. Use your `Square Identity` in the denominator. \(\large 1+\tan^2\theta=?\)

OpenStudy (anonymous):

1+tan^2=sec^2x?

zepdrix (zepdrix):

\[\Large \int\limits\limits\frac{9\left(\frac{1}{3}\sec^2\theta \;d \theta\right)}{1+\tan^2\theta}\qquad=\qquad \int\limits\limits\frac{9\left(\frac{1}{3}\sec^2\theta \;d \theta\right)}{\sec^2\theta}\]Yesss good good.